Mês  QUANTIA
Janeiro 257.655.543
fevereiro 4.567.68
Março 12.356.689
Abril  5.6

Tabela 1: Decimal tamanho pré- definido da coluna em 10ex

y z

t  h j
e  r t

Tabela 2: tamanho fixo

Tabela Decimal

\begin{table}
\centering
\begin{tabular}{|c|r @{.}l|}
\hline
Mês        &  \multicolumn{2}{c|}{QUANTIA} \\   \hline
Janeiro    & 257.655 & 543 \\  \hline
fevereiro  & 4.567   & 68  \\  \hline
Março      & 12.356  & 689 \\  \hline
Abril      & 5       & 6   \\  \hline
\end{tabular}
\caption{Decimal}
\end{table}

Tabela com Tamanho Fixo

\begin{table}
\centering
\begin{tabular}{|p{10ex}|c|c|}
\hline
  tamanho pré-definido
  da coluna em 10ex    & y & z \\ \cline{1-2}
  t                    & h & j \\ \cline{1-1} \cline{3-3}
  e                    & r & t \\
\hline
\end{tabular}
\caption{tamanho fixo}\label{tab:tafix}
\end{table}